OpenStudy (anonymous):

use Simpson rule with 5 function values to find the volume of the solid formed when curve y=e^x is rotated y axis from y=3 to y=5

OpenStudy (anonymous):

i squared y=e^x and it should be Y^2 =e^(2x)

OpenStudy (isaiah.feynman):

Haven't studied this yet.

OpenStudy (anonymous):

ohk nw thanx anyway

OpenStudy (paki):

@mathmale please help @ayeshaafzal221 ...

OpenStudy (paki):

@iambatman

OpenStudy (anonymous):

Do you know the simpsons rule?

OpenStudy (isaiah.feynman):

I'm sure Bruce Wayne knows it. :P

OpenStudy (paki):

@ayeshaafzal221 please respond here, so that you may be able to understand the point...

OpenStudy (anonymous):

i am sorry my internet stuffed up

OpenStudy (anonymous):

ahm yes i know simpson rule i am using

OpenStudy (anonymous):

\[\int\limits_{a}^{b} f(x)dx \approx S _{n} = \frac{ \Delta x }{ 3 }\left[ f(x _{0})+4f(x _{1})+2f(x _{2})+4f(x _{3})+...+2f(x _{n-2})+4f(x _{n-1})+f(x _{n}) \right]\]

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

That ring a bell? It's actually very easy to use

OpenStudy (anonymous):

i know but ahm i am having problem squaring the equation

OpenStudy (anonymous):

y=e^x

OpenStudy (anonymous):

because its rotated y axis so we square the x right?

OpenStudy (anonymous):

Draw it out, it'll help you.

OpenStudy (anonymous):

the drawing is not required

OpenStudy (anonymous):